# GPS-IDS: An Anomaly-based GPS Spoofing Attack Detection Framework for Autonomous Vehicles

### v1

@@ IV-E4 Experiment 4: Tuning of the Detection Threshold $\mathbb{T}$
TABLE III: False Positive (FP) and False Negative (FN) Rates for Different Detection Margins
@@
IV-E5 Experiment 5: Time Series Analysis of the Machine Learning Models
In this experiment, we evaluated the effectiveness of the 4 best-performing models from Experiment 3 (MLP, RF, XGB, and DT) in detecting an attack with respect to time on AV-GPS-Dataset 3. The training phase involved training the models using Case I and Case II methodologies, followed by testing on AV-GPS-Dataset 3. The outcomes of the time series analysis are represented in Fig. 11. As depicted in Fig. 11, the attack was initiated in the $142^{th}$ second and was subsequently detected by the AVT’s EKF algorithm in the $165^{th}$ second (23 seconds delay). In both training cases, all the models succeeded in detecting the attack before the EKF. In Fig. 11(a), MLP was capable of detecting the attack at the $152^{nd}$ second, followed by RF, XGB, and DT at the $158^{th}$ second. Similarly, in Fig. 11(b), the MLP, XGB, and RF identified the attack at the $152^{nd}$ second, while the DT detected it at the $158^{th}$ second. Table IV summarizes the attack detection times for each classifier, revealing that MLP achieved the fastest detection time of 10 seconds in both training methodologies.
@@ IV-E5 Experiment 5: Time Series Analysis of the Machine Learning Models
MLP XGB RF DT EKF
Case I 10 16 16 16 23
Case II 10 10 10 16 23
V Conclusion and Future Work
@@
This paper presents a novel anomaly behavior analysis-based GPS Intrusion Detection framework called the GPS-IDS that detects abnormal GPS navigation in AVs. The approach uses physics-based vehicle modeling to represent the behavior of the vehicle. A modified dynamic bicycle model is utilized to capture the normal behavior of the vehicle and machine learning techniques are used to detect the anomalous behavior. Moreover, a novel dataset family called the AV-GPS-Dataset with over 69,000 instances of real-world autonomous vehicle normal data and GPS spoofing attack data, is introduced in this paper. The performance of the proposed GPS-IDS approach is evaluated using the AV-GPS-Dataset, and the experimental results affirm its high detection rate of GPS spoofing attacks. Additionally, it is validated that the proposed approach exhibits faster detection times in comparison to the EKF-based detection algorithm. In contrast to the existing segregated intrusion detection techniques that concentrate only on individual sensor data, our approach considers the overall behavior of the system, providing a more comprehensive approach to detect GPS attacks. We argue that using GPS-IDS in conjunction with sensor-specific intrusion detection systems can provide a powerful defense mechanism against GPS spoofing attacks in autonomous vehicles.
@@
Moving forward, there are several promising research avenues to extend the contributions of this work. The GPS-IDS algorithm can be refined by incorporating additional operational parameters in the Autonomous Vehicle Behavior Model, such as integrating accurate modeling of multiple sensors and more robust control algorithms. Further research can be done to explore the effectiveness of integrating GPS-IDS with existing sensor-specific detection systems to enhance the overall security of autonomous vehicles. Additionally, the feasibility of real-world deployment scenarios, including cloud environment computing, onboard deployment on resource-constrained edge devices, or hybrid computation strategies that balance cloud and onboard processing can be investigated to validate the scalability and practicality of the GPS-IDS approach. Finally, the development of a digital twin for the autonomous vehicle behavior model emerges as a promising direction, providing a virtual environment for comprehensive testing and simulation.

### v2

@@ V-D4 Experiment 4- Tuning of the Detection Threshold $\mathbb{T}$
TABLE III: False Positive (FP) and False Negative (FN) Rates for Different Detection Margins in Real Datasets
@@
V-D5 Experiment 5- Time Series Analysis of the Machine Learning Models
In this experiment, we evaluated the effectiveness of the 4 best-performing models from Experiment 3 (MLP, RF, XGB, and DT) in detecting an attack with respect to time on AV-GPS-Dataset 3. The training phase involved training the models using Case I and Case II methodologies, followed by testing on AV-GPS-Dataset 3. The outcomes of the time series analysis are represented in Fig. 10. As depicted in Fig. 10, the attack was initiated in the $142^{th}$ second and was subsequently detected by the AVT’s EKF algorithm in the $165^{th}$ second (23 seconds delay). In both training cases, all the models succeeded in detecting the attack before the EKF. In Fig. 10(a), MLP was capable of detecting the attack at the $152^{nd}$ second, followed by RF, XGB, and DT at the $158^{th}$ second. Similarly, in Fig. 10(b), the MLP, XGB, and RF identified the attack at the $152^{nd}$ second, while the DT detected it at the $158^{th}$ second. Table IV summarizes the attack detection times for each classifier, revealing that MLP achieved the fastest detection time of 10 seconds in both training methodologies and showcasing a $\approx$ 56.5% improvement in detection time compared to GPS/INS-based detection method using EKF.
@@ V-D5 Experiment 5- Time Series Analysis of the Machine Learning Models
MLP XGB RF DT EKF Max. Improve
Case I 10 16 16 16 23 $\approx$ 56.52%
Case II 10 10 10 16 23 $\approx$ 56.52%
V-D6 Experiment 6- Performance Analysis of the Machine Learning Models with Different Simulated Training Sets
To ensure the generalization of the GPS-IDS approach, this experiment evaluated GPS-IDS on the simulation-generated AV-GPS-Datasets 4—6 using different machine learning models. Six models were chosen: RF, XGB, SVC, MLP, AdaBoost, and GB. Similar to Experiment 3, supervised learning was utilized and stratified 5-fold sampling was employed, where 4 folds (80%) were used for training, and the remaining fold (20%) was used for testing. Three cases were considered for training, Case IV: Train on 80% of AV-GPS-Dataset 4, and test on 20% of AV-GPS-Dataset 4, AV-GPS-Dataset 5, and AV-GPS-Dataset 6; Case V: Train on 80% of AV-GPS-Dataset 5, and test on AV-GPS-Dataset 4, 20% of AV-GPS-Dataset 5 and AV-GPS-Dataset 6; Case VI: Train on 80% of AV-GPS-Dataset 6, and test on AV-GPS-Dataset 4, AV-GPS-Dataset 5, and 20% of AV-GPS-Dataset 6. For all three cases, the performance of the machine learning models is measured in terms of Accuracy, Precision, Recall, and F1-score. For each model, we used the same tuned hyperparameters utilized as Experiment 3. The performance of machine learning algorithms on different simulation generated datasets is presented in Table V. The results indicate that the performance varies across different datasets as the training set changes. MLP consistently performs with F1 scores above 93% in all three cases, achieving the highest average F1 score of 97.1%. Comparing Table II and Table V, we can see that the GPS-IDS approach demonstrates an overall better performance on the simulation-based datasets.
TABLE V: Performance of the Machine Learning Classification Models on AV-GPS-Datasets 4—6 (Simulation-generated Data)
AV-GPS-Datasets 4—6 Metrics RF XGB SVC MLP Adaboost GB
Case IV Trained with 80% of Dataset 4 Test on 20% of Dataset 4 Accuracy 0.8724 0.9495 0.9091 0.9471 0.9283 0.9285
Precision 0.8955 0.9541 0.9093 0.9507 0.9294 0.9295
Recall 0.8724 0.9495 0.9091 0.9471 0.9283 0.9285
F1 Score 0.8705 0.9494 0.9091 0.9470 0.9283 0.9284
Test on Dataset 5 Accuracy 0.9494 0.9494 0.9523 0.9790 0.9486 0.9068
Precision 0.9541 0.9541 0.9564 0.9798 0.9534 0.9215
Recall 0.9494 0.9494 0.9523 0.9790 0.9486 0.9068
F1 Score 0.9493 0.9493 0.9522 0.9789 0.9485 0.9060
Test on Dataset 6 Accuracy 0.9492 0.9492 0.9527 0.9518 0.9500 0.9151
Precision 0.9539 0.9539 0.9568 0.9560 0.9545 0.9275
Recall 0.9492 0.9492 0.9527 0.9518 0.9500 0.9151
F1 Score 0.9491 0.9491 0.9526 0.9517 0.9499 0.9145
Average F1 Score in Case I 0.9230 0.9493 0.9380 0.9592 0.9422 0.9163
Case V Trained with 80% of Dataset 5 Test on Dataset 4 Accuracy 0.9409 0.9454 0.9590 0.9619 0.9040 0.9835
Precision 0.9471 0.9493 0.9621 0.9646 0.9040 0.9837
Recall 0.9409 0.9454 0.9590 0.9619 0.9040 0.9835
F1 Score 0.9407 0.9453 0.9589 0.9618 0.9040 0.9835
Test on 20% of Dataset 5 Accuracy 0.8591 0.9397 0.9492 0.9379 0.8911 0.9285
Precision 0.8760 0.9427 0.9538 0.9408 0.8981 0.9295
Recall 0.8591 0.9397 0.9492 0.9379 0.8911 0.9285
F1 Score 0.8575 0.9396 0.9491 0.9378 0.8906 0.9284
Test on Dataset 6 Accuracy 0.9461 0.9492 0.9987 0.9996 0.9635 0.9151
Precision 0.9513 0.9539 0.9987 0.9996 0.9660 0.9275
Recall 0.9461 0.9492 0.9987 0.9996 0.9635 0.9151
F1 Score 0.9459 0.9491 0.9987 0.9996 0.9635 0.9145
Average F1 Score in Case II 0.9147 0.9450 0.9689 0.9664 0.9194 0.9421
Case VI Trained with 80% of Dataset 6 Test on Dataset 4 Accuracy 0.9429 0.9400 0.9599 0.9721 0.9409 0.9511
Precision 0.9488 0.9430 0.9628 0.9733 0.9432 0.9554
Recall 0.9429 0.9400 0.9599 0.9721 0.9409 0.9511
F1 Score 0.9427 0.9399 0.9598 0.9721 0.9409 0.9511
Test on Dataset 5 Accuracy 0.9986 0.9494 0.9998 0.9997 0.9995 0.9962
Precision 0.9986 0.9541 0.9998 0.9997 0.9995 0.9963
Recall 0.9986 0.9494 0.9998 0.9997 0.9995 0.9962
F1 Score 0.9986 0.9493 0.9998 0.9997 0.9995 0.9962
Test on 20% of Dataset 6 Accuracy 0.8268 0.9090 0.9522 0.9910 0.9349 0.9107
Precision 0.8602 0.9229 0.9564 0.9912 0.9360 0.9214
Recall 0.8268 0.9090 0.9522 0.9910 0.9349 0.9107
F1 Score 0.8227 0.9082 0.9521 0.9910 0.9348 0.9101
Average F1 Score in Case III 0.9213 0.9325 0.9706 0.9876 0.9584 0.9525
Average F1 Score of all three cases 0.9197 0.9423 0.9592 0.9711 0.9400 0.9370
V-D7 Experiment 7- Tuning of the Detection Threshold $\mathbb{T}$ on Simulated Data
Similar to Experiment 4, this experiment assessed the sensitivity of the GPS-IDS detection threshold $\mathbb{T}$. We varied threshold values to evaluate their impacts on GPS-IDS performance using the simulation-generated datasets. The machine learning model used for this experiment was the MLP, given its consistent best performance in the previous experiments. The resulting data were plotted to determine a suitable detection margin, as illustrated in Fig. 11 By comparing the probability score distributions for normal and attack data, we again find that selecting a detection threshold $\mathbb{T}$ within a range of 0.4 to 0.5 shows the best performance, resulting in misclassification of only 295 instances of normal data and no instance of attack data and leading to a False Positive Rate of 0.0076 or 0.76% and False Negative Rate of 0%. Table. VI summarizes the different False Positive and False Negative rates for different margins in simulated data.
TABLE VI: False Positive (FP) and False Negative (FN) Rates for Different Detection Margins in Simulated Datasets
0.4 - 0.5 295 0 0.0076 0.0
0.3 - 0.5 764 0 0.0198 0.0
0.4 - 0.6 295 200 0.0076 0.0052
0.3 - 0.6 764 200 0.0198 0.0052
Figure 11: Comparison of probability score distribution for normal and attack data in simulations
VI Conclusion and Future Work
@@
This paper presents a novel anomaly behavior analysis-based GPS Intrusion Detection framework called the GPS-IDS that detects abnormal GPS navigation in AVs. The approach uses physics-based vehicle modeling to represent the behavior of the vehicle. A modified dynamic bicycle model is utilized to capture the normal behavior of the vehicle and machine learning techniques are used to detect the anomalous behavior. Moreover, a novel dataset family called the AV-GPS-Datasets with both real-world and simulated data is introduced in this paper. The performance of the proposed GPS-IDS approach is evaluated using the AV-GPS-Dataset, and the experimental results affirm its high detection rate of GPS spoofing attacks. Additionally, it is validated that the proposed approach exhibits faster detection times in comparison to the EKF-based detection algorithm in real data of AV-GPS-Datasets. In contrast to the existing segregated intrusion detection techniques that concentrate only on individual sensor data, our approach considers the overall behavior of the system, providing a more comprehensive approach to detect GPS attacks. We argue that using GPS-IDS in conjunction with sensor-specific intrusion detection systems can provide a powerful defense mechanism against GPS spoofing attacks in autonomous vehicles. To ensure the generalization of our approach, we conducted simulation-based experiments to show that our suggested system can perform equally well in a more complex urban scenario with multiple vehicles under attack.
@@
Moving forward, there are several promising research avenues to extend this work. The GPS-IDS algorithm can be refined by incorporating additional operational parameters in the Autonomous Vehicle Behavior Model, such as integrating accurate modeling of multiple sensors and more robust control algorithms. Further research can be done to explore the effectiveness of integrating GPS-IDS with existing sensor-specific detection systems to enhance the overall security of AVs. Additionally, GPS-IDS can be improved to detect more stealthy GPS attacks, such as dual-band spoofing and spoofing when aligned with road shapes [68, 69] against AVs. Furthermore, the feasibility of real-world deployment scenarios, including cloud environment computing, onboard deployment on resource-constrained edge devices, or hybrid computation strategies that balance cloud and onboard processing can be investigated to validate the scalability and practicality of the GPS-IDS approach. Finally, the development of a digital twin for the Autonomous Vehicle Behavior Model emerges as a promising direction, providing a virtual environment for comprehensive testing and simulation.